# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Abstract
The Fourier Transform (FT) is a fundamental tool that permeates modern science and technology. While chemistry undergraduates encounter the FT as early as second year, their courses often only mention it in passing because computers frequently perform it automatically behind the scenes. Although this automation enables students to focus on ‘the chemistry’, students miss out on an opportunity to understand and use one of the most powerful tools in the scientific arsenal capable of revealing how time-dependent signals encode chemical structure. Although many educational resources introduce chemists to the FT, they often require familiarity with sophisticated mathematical and computational concepts. Here, we present a series of three self-contained, Python-based laboratory activities for undergraduates to understand the FT and apply it to analyze audio signals, an infrared (IR) spectroscopy interferogram, and a nuclear magnetic resonance (NMR) free induction decay (FID). In these activities, students observe how the FT reveals and quantifies the contribution of each frequency present in a temporal signal and how decay timescales dictate signal broadening. Our activities empower students with the tools to transform their own temporal datasets (e.g., FID) to a frequency spectrum. To ensure accessibility of the activities and lower the barrier to implementation, we utilize Google Colab’s open-source, cloud-based platform to run Jupyter notebooks. We also offer a pre-laboratory activity that introduces students to the basics of Python and the Colab platform, and reviews the math and programming skills needed to complete the lab activities. These lab activities help students build a qualitative, quantitative, and practical understanding of the FT.
@@ I Introduction
We have tested these activities as part of a three-day ‘Math and Physical Chemistry BootCamp’ for first-year physical chemistry graduate students and in two upper-division physical chemistry lab courses, CHEM 4250 at Trinity University and CHM 317L at Chatham University. As written, the activities require two three-hour laboratory sessions. However, we have designed these activities to be modular and also provide a one-hour streamlined version that an instructor could incorporate into a single lecture session (Supporting Information).
@@ II Theoretical Background
Here, we outline the background mathematics students and instructors need to understand the FT and complete the lab activities. For more detailed discussions on these topics, we refer the reader to Refs. Mark Newman, 2013; Pres et al., 1992; Griffiths, 1978. We then provide an overview of NMR and IR spectroscopy theory and refer the reader to Refs. Akitt, 1983; Günther, 1973 for more detailed explanations. We additionally provide detailed derivations in the Supporting Information.
@@ II.1 Fourier Transforms
Although we commonly employ the FT to decompose a time-domain signal, $f(t)$, into its frequency components, $\hat{f}(\omega)$, it is far more general. Fundamentally, the FT connects ‘conjugate variables’, e.g., time and frequency, or position and momentum. Additionally, one can apply the FT to periodic functions that repeat over a finite domain or to aperiodic functions that decay over an infinite domain. We illustrate both cases in detail in this section.
@@ II.1.1 FT over a finite domain
$f(t)=\sum_{n=0}^{\infty}a_{n}\cos\left(\omega_{n}t\right)+\sum_{n=1}^{\infty}b_{n}\sin\left(\omega_{n}t\right),$ (1)
@@
where $\omega_{n}=\pi n/T$ is the angular frequency of the sinusoid, and the coefficients $a_{n}$ and $b_{n}$ contribute to the Fourier coefficients that compose the spectrum. The expansion in Eq. (1) is the Fourier series of $f(t)$. To clearly connect the FT on finite and infinite domains, we recast Eq. (1) as
$f(t)=\sum_{n=-\infty}^{\infty}c_{n}e^{-i\omega_{n}t},$ (2)
@@
where $i=\sqrt{-1}$ is the complex unit, $e^{ix}=\cos(x)+i\sin(x)$ is Euler’s identity for decomposing a complex exponential into sinusoids, and the (possibly complex) coefficients $\{c_{n}\}$ are given by the integral
$c_{n}=\frac{1}{T}\int_{-T}^{T}f(t)e^{-i\omega_{n}t}\,{\rm d}t.$ (3)
@@
The collection of Fourier coefficients $\{c_{n}\}$ are the weights of specific frequency contributions in the original signal $f(t)$ and constitute the frequency spectrum. Together, Eqs. (3) and (2) suggest that the FT is invertible, meaning that not only can one compute the spectrum of $f(t)$ using Eq. (3), one can then subsequently ‘undo’ this transformation using Eq. (2) to obtain the original temporal signal.
@@
Qualitatively, one interprets the coefficient $c_{n}$ as the amount of the $n^{\rm th}$ frequency contained in the temporal signal $f(t)$. The fundamental structure that quantifies this amount is the ‘orthogonality’ of trigonometric functions. Mathematically, if two nonzero functions $f(t)\neq g(t)$ are orthogonal, then their product integrates to zero over their defined interval, $\int_{-T}^{T}f(t)g(t)\,{\rm d}t=0$. However, if this integral evaluates to a nonzero value, then this value tells you ‘how much’ these two functions overlap. With this in mind, we can answer questions Q1 and Q2 from the beginning of this section.
@@
In Fig. 2, we demonstrate the application of the finite-domain FT by considering the periodic temporal signal $f(t)=5\cos(t)+10\cos(3t)$ on $-\pi\leq t<\pi$ (Fig. 2A). The function $f$ is a combination of waves with frequencies $\omega=1$ Hz and $\omega=3$ Hz. Therefore, we expect to see in $c_{n}$, the FT of $f(t)$, a peak at each frequency represented in $f(t)$ with heights defined by the amplitudes of each sinusoid, as shown in Fig. 2B. To illustrate the origin of each peak in Fig. 2B, we graph the integrand at each of three trial frequencies ($\omega=1,2,3$ Hz). For example, in Fig. 2C, we consider $\omega_{1}=1$ Hz and plot the product $f(t)e^{-it}$. The shaded area under this curve corresponds to how much these two functions overlap and the numerical value of this area is equivalent to the integral in Eq. (3). As expected, the shaded area is nonzero because there is overlap between our input signal $f(t)$ and $e^{-it}=\cos(t)-i\sin(t)$ since they both contain $\cos(t)$, and thus we see a peak at $c_{1}$ in Fig. 2. For similar reasons, we would expect that the area under $f(t)e^{-3it}$ would be nonzero (Fig. 2E), since both $f(t)e^{-3it}$ contain $\cos(3t)$. The peak at $c_{3}$ is larger than $c_{1}$ because the amplitude corresponding to $\omega_{3}$ is greater than the amplitude corresponding to $\omega_{1}$ and thus there is ‘more of’ $\omega_{3}$ than $\omega_{1}$ in the input signal $f(t)$. Unlike both of these cases, in Fig. 2D we observe that the area under $f(t)e^{-2it}$ is zero. This is because $f(t)$ does not contain any components with a frequency of $\omega_{2}=2$ Hz and therefore the $e^{-2it}$ and $f(t)$ are orthogonal.
@@
Figure 2: Illustration of the FT over a finite domain (see Sec. II.1.1). (A) Temporal signal created from the addition of two cosines with two distinct frequencies. (B) Frequency spectrum obtained by computing the Fourier coefficients $c_{n}$ via Eq. (3). (C)-(E) Demonstrating the orthogonality of the trig functions by multiplying the temporal signal $f(t)$ by the complex exponential $e^{-i\omega_{n}t}$, where $\omega_{n}=2\pi n/T$ Hz with $T=2\pi$ seconds. The shaded area corresponds to the integral in Eq. (3), thus enabling one to visually conclude that $c_{1}<c_{3}$ and that $c_{2}=0.$
II.1.2 FT over an infinite domain
@@
In the above section, the underlying assumption is that the signal $f(t)$ is periodic with period $2T$, i.e., the same function repeats identically on all intervals $f(t+2T)=f(t)$. Here we show how to extend this approach to time-series data that are not indefinitely periodic, like the free induction decay measurements obtained in a proton NMR experiment. The trick is to view the signal as having an infinite period length. By applying the limit definition of the integral from calculusStewart (2020); Pons (2014) in conjunction with the limit as $T\rightarrow\infty$ (see Supporting Information) one can manipulate Eq. (3) into
@@
the frequency spectrum of our temporal, aperiodic signal. Similar to the discussion in Sec. II.1.1, one can also ‘undo’ Eq. 4 and return back to the original signal $f(t)$ using the inverse FT
@@
For completeness, we note that the above integrals exist provided that both $\lim_{t\rightarrow\pm\infty}f(t)=0$ and $\lim_{\omega\rightarrow\pm\infty}\hat{f}(\omega)=0$, respectively, and are absolutely integrable (i.e., $\int_{-\infty}^{\infty}|f(t)|\,{\rm d}t<\infty$ and $\int_{-\infty}^{\infty}|\hat{f}(\omega)|\,{\rm d}\omega<\infty$). Generally, spectroscopic measurements (e.g., FIDs, interferograms) meet these requirements.
@@ II.1.3 Practical implementation of the FT
In practice, the measured time-domain signal $f(t)$ is sampled over finite, nonnegative values of time $0\leq t<\infty$ with finite sampling intervals $\Delta t$. Thus, we consider the discrete analogue of Eq. (4) termed the discrete Fourier transform (DFT). Mathematically, the DFT is given by
$\hat{f}(\omega_{n})=\sum_{k=0}^{N-1}f(t_{k})e^{-2\pi i\omega_{n}t_{k}},$ (6)
where $k$ is an integer that takes values from 00 to $N-1$, and $N$ is the number of measurements in the dataset.Pres et al. (1992)
Although one can implement the FT directly Eq. (6), this approach is often inefficient. Specifically, calculating the DFT directly at each frequency $\omega_{n}$ requires evaluating the sum in Eq. (6) over all values of time $t_{k}$. Thus, evaluating $\hat{f}$ computationally requires two nested for-loops: one running over $\omega_{n}$ and the other over $t_{k}$. Thus, when one has $N$ values of $t_{k}$ and $N$ values of $\omega_{n}$, the total number of operations scales as $N^{2}$, quantifying the cost of the direct algorithm. In computer science language, we say that the computational complexity of the DFT algorithm scales quadratically with the size of the time-domain dataset, i.e., $\mathcal{O}(N^{2})$ scaling.Pres et al. (1992) We illustrate this scaling issue in Section III.2.4 to motivate the introduction of “the most important numerical algorithm of our lifetime”Strang (1994): the fast Fourier transform (FFT),Pres et al. (1992) which reduces the scaling to $\mathcal{O}(N\log_{2}N)$,Pres et al. (1992) rendering the FFT as a significantly more efficient practical algorithm to computing the Eq. (6).
@@ II.2 FT-NMR
NMR spectroscopy is a topic familiar to all undergraduate chemistry students, but the role of the FT in data collection is rarely a point of emphasis. While, in principle, NMR spectroscopy can be conducted in continuous wave mode with a fixed magnetic field and sequentially collecting data at each frequency, almost all modern NMR spectrometers use the FT mode.
@@
For completeness, we review the basics of NMR spectroscopy to illustrate to the interested student how NMR offers a means to distinguish chemical environments. NMR spectroscopy exploits interactions between atomic nuclei (e.g., 1H and 13C) and an external magnetic field ${\bf B}$. In the presence of a magnetic field aligned in the $x$-direction with strength $B_{x}=|{\bf B}|$, the nuclear spins of the chemical sample either align or anti-align with the direction of the field, leading to an energy difference between those aligned and anti-aligned given by $\Delta E=\tfrac{\mu B_{x}}{I},$ where $\mu$ and $I=\pm 1/2$ are the magnetic dipole moment (values found at Ref. Stone, 2005) and spin for 1H nuclei, respectively. In FT-NMR, the sample is subjected to a radiofrequency (RF) pulse that perturbs the external magnetic field. The energy of this RF pulse is tuned to the field-dependent frequency, $f_{0}$, of the atomic nucleus one is targeting (e.g., 1H, 13C), given by its gyromagnetic ratio, $\gamma$, via $f_{0}=\gamma B_{x}$.Akitt (1983); Günther (1973) This perturbation results in a change in the orientation of the nuclear spins within the bulk sample. Following this RF pulse, the nuclear spins relax back to their lower energy configuration (ground state) under the constant field, $B_{x}$. As these spins relax, they emit a RF signal, and this causes a measurable change in the induced voltage in the NMR detector coil; we call this signal the free induction decay (FID). The FT of the FID results in the familiar frequency domain spectrum that students learn in organic chemistry.
@@
The spectrum we obtain in 1H-NMR experiments provides insight into the molecular structure of a sample. After applying the FT to a temporal signal, a typical spectrum shows intensity as a function of frequency $\omega$ in units of Hertz. However, the convention is to convert this frequency in Hertz to a chemical shift ($\delta$) with units of parts-per-million (ppm) which reports on the change in frequency relative to the reference frequency of tetramethyl silane (TMS), $\omega_{\rm TMS}$, and the operational frequency of the spectrometer, $\omega_{\rm spec}$.Akitt (1983) Mathematically, $\delta=\tfrac{\omega-\omega_{\rm TMS}}{\omega_{\rm spec}}\times 10^{6}$. Importantly, this conversion yields a field-independent measurement and implies that the chemical shift does not change from instrument to instrument.
@@
Students are likely familiar with using features of a FT-NMR spectrum (e.g., chemical shifts, $J$-couplings, peak integrations), to characterize the structure of a molecule. It turns out that all of this same information is encoded into the FID, but in a more cryptic way (see Secs. II.1.1 and II.1.2). To see this, we recall that the FID signal encodes the frequency each proton nucleus emits in response to the initial RF pulse. Because this response depends on the nuclei’s chemical and electrical environments, the emitted RF signal is a sinusoid composed of frequencies unique to each nucleus’s environment. These signals combine to form the FID signal that the spectrometer measures at the detector. In addition, this FID signal decays to zero as the nuclear spins relax and approach equilibrium, with the relaxation being dictated by the chemical environment. The timescale of this relaxation is given by the spin-spin relaxation constant ($T_{2}$), which helps determine the height and width of each peak in the spectrum.Akitt (1983) Because the FID signal is a combination of sinusoids with various frequencies and the signal decays to zero, it satisfies the requirements for the FT, as discussed in Sec. II.1.

### v2

@@ Abstract
The Fourier Transform (FT) is a fundamental tool that permeates modern science and technology. While chemistry undergraduates encounter the FT as early as the second year, their courses often only mention it in passing because computers frequently perform it automatically behind the scenes. Although this automation enables students to focus on ‘the chemistry’, students miss out on an opportunity to understand and use one of the most powerful tools in the scientific arsenal capable of revealing how time-dependent signals encode chemical structure. Although many educational resources introduce chemists to the FT, they often require familiarity with sophisticated mathematical and computational concepts. Here, we present a series of three self-contained, Python-based laboratory activities designed for undergraduates to understand the FT and apply it to analyze audio signals, an infrared (IR) spectroscopy interferogram, and a nuclear magnetic resonance (NMR) free induction decay (FID). In these activities, students observe how the FT reveals and quantifies the contribution of each frequency present in a temporal signal and how decay timescales dictate signal broadening. Our activities empower students with the tools to transform their own temporal datasets (e.g., FID) to a frequency spectrum. To ensure accessibility of the activities and lower the barrier to implementation, we utilize Google Colab’s open-source, cloud-based platform to run Jupyter notebooks. We also offer a pre-laboratory activity that introduces students to the basics of Python and the Colab platform, and reviews the math and programming skills needed to complete the lab activities. These lab activities help students build a qualitative, quantitative, and practical understanding of the FT.
@@ I Introduction
We have tested these activities as part of a three-day ‘Math and Physical Chemistry BootCamp’ for first-year physical chemistry graduate students and in two upper-division physical chemistry lab courses, CHEM 4250 at Trinity University and CHM 317L at Chatham University. As written, the activities require two three-hour laboratory sessions. However, we have designed these activities to be modular and also provide a one-hour streamlined version that an instructor could incorporate into a single lecture session (see Streamlined.ipynb in the Supporting Information).
@@ II Theoretical Background
We designed the lab activities in Sec. III to be accessible to junior and senior chemistry undergraduate students. Specifically, we do not assume that students have any prior experience with the FT, and we have found that students can complete the activities with no prerequisites other than qualitative familiarity with spectroscopy obtained in undergraduate organic chemistry and familiarity with the tools of calculus. Nevertheless, while a detailed theoretical understanding of the FT is not required for students, we provide a brief, pedagogical overview of the theory of the FT Mark Newman (2013); Pres et al. (1992); Griffiths (1978) and its applications to NMRGünther (1973); Akitt (1983) and IRGriffiths and De Haseth (2007) spectroscopy for the benefit of the instructor. These can also be offered to interested students after they complete the activities for further discussion and enrichment. We describe the design and implementation of the lab activities in Sec. III.
@@ II.1 Fourier Transforms
Although we commonly employ the FT to decompose a time-domain signal, $f(t)$, into its frequency components, $\hat{f}(\omega)$, its applicability is far more general. Fundamentally, the FT connects ‘conjugate variables’, e.g., time and frequency, or position and momentum. Additionally, one can apply the FT to periodic functions that repeat over a finite domain (Fourier series) or to aperiodic functions that decay over an infinite domain (Fourier transform). We illustrate both cases in detail in this section.
@@ II.1.1 Fourier Series
$f(t)=\sum_{n=0}^{\infty}a_{n}\cos\left(\omega_{n}t\right)+\sum_{n=1}^{\infty}b_{n}\sin\left(\omega_{n}t\right)$ (1)
@@
where $\omega_{n}=\pi n/T$ is the angular frequency of the sinusoid, and the coefficients $a_{n}$ and $b_{n}$ contribute to the Fourier coefficients that compose the spectrum. The expansion in Eq. (1) is the Fourier series of $f(t)$. To connect the Fourier series for periodic functions and the FT for aperiodic functions, one can rewrite Eq. (1) into the more compact form
$f(t)=\sum_{n=-\infty}^{\infty}\hat{f}_{n}e^{i\omega_{n}t},$ (2)
@@
where $i=\sqrt{-1}$ is the complex unit, $e^{ix}=\cos(x)+i\sin(x)$ is Euler’s identity for decomposing a complex exponential into sinusoids, and the (possibly complex) Fourier coefficients $\{\hat{f}_{n}\}$ are given by the integral
$\hat{f}_{n}=\frac{1}{T}\int_{-T}^{T}f(t)e^{-i\omega_{n}t}\,{\rm d}t.$ (3)
@@
We include this derivation in the Appendix VIII.1 for the interested reader. Here, we decorate $f_{n}$ with a ‘hat’, $\hat{f}_{n}$, to denote the FT of the signal $f(t)$. Together, Eqs. (3) and (2) suggest that the FT is invertible, meaning that not only can one compute the spectrum of $f(t)$ using Eq. (3), one can then subsequently ‘undo’ this transformation using Eq. (2) to obtain the original temporal signal.
@@
Qualitatively, one interprets the coefficient $\hat{f}_{n}$ as the amount of the $n^{\rm th}$ frequency contained in the temporal signal $f(t)$. The mathematical concept that clarifies the meaning of the Fourier integral (Eq. (3)) is the ’orthogonality’ of trigonometric functions. Mathematically, if two nonzero functions $f(t)\neq g(t)$ are orthogonal, their product integrates to zero over the interval $-T\leq t\leq T$, i.e., $\int_{-T}^{T}f(t)g(t)\,{\rm d}t=0$. However, if this integral evaluates to a nonzero value, then this value tells you ‘how much’ these two functions overlap. With this in mind, we can answer questions Q1 and Q2 from the beginning of this section.
@@
In Fig. 2, we demonstrate how the orthogonality concept helps us interpret the Fourier series by considering the periodic temporal signal $f(t)=5\cos(t)+10\cos(3t)$ on $-\pi\leq t<\pi$ (Fig. 2A). The function $f$ is a combination of waves with frequencies $\omega=1$ s-1 and $\omega=3$ s-1. Therefore, we expect to see in $\hat{f}_{n}$, the FT of $f(t)$, a peak at each frequency represented in $f(t)$ with heights defined by the amplitudes of each sinusoid, as shown in Fig. 2B. To illustrate the origin of each peak in Fig. 2B, we graph the integrand at each of three trial frequencies ($\omega=1,2,3$ s-1). For example, in Fig. 2C, we consider $\omega_{1}=1$ s-1 and plot the product ${\rm Re}[f(t)e^{-it}]$. The shaded area under this curve corresponds to how much these two functions overlap and the numerical value of this area is equivalent to the integral in Eq. (3). As expected, the shaded area is nonzero because there is overlap between our input signal $f(t)$ and $e^{-it}=\cos(t)-i\sin(t)$ since they both contain $\cos(t)$, and thus we see a peak at $f_{1}$ in Fig. 2. For similar reasons, we would expect that the area under ${\rm Re}[f(t)e^{-3it}]$ would be nonzero (Fig. 2E), since both ${\rm Re}[f(t)e^{-3it}]$ contain $\cos(3t)$. The peak at $f_{3}$ is larger than $f_{1}$ because the amplitude corresponding to $\omega_{3}$ is greater than the amplitude corresponding to $\omega_{1}$ and thus there is ‘more of’ $\omega_{3}$ than $\omega_{1}$ in the input signal $f(t)$. Unlike both of these cases, in Fig. 2D we observe that the area under ${\rm Re}[f(t)e^{-2it}]$ is zero. This is because $f(t)$ does not contain any components with a frequency of $\omega_{2}=2$ s-1 and therefore the $e^{-2it}$ and $f(t)$ are orthogonal. Thus, this example shows that the collection of Fourier coefficients $\{\hat{f}_{n}\}$ are the weights of the specific frequency contributions, and the plot of $\hat{f}_{n}$ vs. $\omega_{n}$ is the frequency spectrum.
@@
Figure 2: Illustration of the Fourier series (see Sec. II.1.1). (A) Temporal signal created from the addition of two cosines with two distinct frequencies. (B) Frequency spectrum obtained by computing the Fourier coefficients $\hat{f}_{n}$ via Eq. (3). (C)-(E) Demonstrating the orthogonality of the trigonometric functions by considering the real component of the product of the temporal signal $f(t)$ and the complex exponential $e^{-i\omega_{n}t}$, where $\omega_{n}=2\pi n/T$ with $T=\pi$ seconds. Note that ${\rm Im}[e^{-i\omega_{n}t}f(t)]=0$ in this example. The shaded area corresponds to the integral in Eq. (3), thus enabling one to visually conclude that $f_{1}<f_{3}$ and that $f_{2}=0.$
II.1.2 The Fourier Transform
@@
In the above section, the underlying assumption is that the signal $f(t)$ is periodic with period $2T$, i.e., the same function repeats identically on all intervals $f(t+2T)=f(t)$. Here we show how to extend this approach to time-series data that are not indefinitely periodic, like the free induction decay measurements obtained in a proton NMR experiment. The trick is to view the signal as having an infinite period length. By applying the limit definition of the integral from calculusStewart (2020); Pons (2014) in conjunction with the limit as $T\rightarrow\infty$ (see Appendix VIII.2) one can manipulate Eq. (3) into
@@
the frequency spectrum of our temporal, aperiodic signal. As we mentioned in Sec. II.1.1, we use the ‘hat’, $\hat{f}$, to denote the FT of the signal $f$. Similar to the discussion in Sec. II.1.1, one can also ‘undo’ Eq. 4 and return back to the original signal $f(t)$ using the inverse FT [^1] For completeness, we note that the above integrals exist provided that $\lim_{t\rightarrow\pm\infty}f(t)=0$ and $\lim_{\omega\rightarrow\pm\infty}\hat{f}(\omega)=0$, respectively, and $f$ and $\hat{f}$ are absolutely integrable (i.e., $\int_{-\infty}^{\infty}|f(t)|\,{\rm d}t<\infty$ and $\int_{-\infty}^{\infty}|\hat{f}(\omega)|\,{\rm d}\omega<\infty$). Generally, spectroscopic measurements (e.g., FIDs, interferograms) meet these requirements.
@@ II.1.3 Practical implementation of the FT
In practice, the measured time-domain signal $f(t)$ is sampled over finite, nonnegative values of time $0\leq t<\infty$ with finite sampling intervals $\Delta t$, and one computes the FT $\hat{f}$ at discrete values of $\omega$. Thus, we consider the discrete analogue of Eq. (4) termed the discrete Fourier transform (DFT). Mathematically, the DFT is given by
$\hat{f}(\omega_{n})=\sum_{k=0}^{N-1}f(t_{k})e^{-i\omega_{n}t_{k}},$ (6)
where $k$ is an integer time index that takes values from 00 to $N-1$, and $N$ is the number of measurements in the dataset, and $n$ is an integer frequency index.Pres et al. (1992)
Although one can implement the FT in Eq. (6) directly, this approach is often inefficient because calculating the DFT directly at each frequency $\omega_{n}$ requires evaluating the sum in Eq. (6) over all values of time $t_{k}$. That is, evaluating $\hat{f}$ for all frequencies $\omega_{n}$ computationally requires two nested for-loops: one running over $\omega_{n}$ and the other over $t_{k}$. Thus, when one has $N$ values of $t_{k}$ and $N$ values of $\omega_{n}$, the total number of operations scales as $N^{2}$, quantifying the cost of the direct algorithm. In computer science language, we say that the computational complexity of the DFT algorithm scales quadratically with the size of the time-domain dataset, i.e., $\mathcal{O}(N^{2})$ scaling.Pres et al. (1992) We illustrate this scaling issue in Sec. III.2.4 to motivate the introduction of “the most important numerical algorithm of our lifetime”Strang (1994): the fast Fourier transform (FFT),Pres et al. (1992) which reduces the scaling to $\mathcal{O}(N\log_{2}N)$,Pres et al. (1992) rendering the FFT as a significantly more efficient and practical algorithm for computing Eq. (6). We refer readers interested in the FFT to Ref. Pres et al., 1992.
@@ II.2 FT-NMR
NMR spectroscopy is a topic familiar to all undergraduate chemistry students, but the role of the FT in data collection is rarely a point of emphasis. While, in principle, one can perform NMR spectroscopy using a continuous wave mode with a fixed magnetic field and sequentially collecting data at each frequency (frequency-sweep method) or with a fixed frequency and a variable magnetic field (field-sweep method),Akitt (1983); Günther (1973) almost all modern NMR spectrometers use the FT mode that we describe below.
@@
For completeness, we review the basics of NMR spectroscopy for spin 1/2 nuclei to illustrate to the interested student how NMR offers a means to distinguish chemical environments. NMR spectroscopy exploits interactions between atomic nuclei (e.g., 1H and 13C) and an external magnetic field ${\bf B}$. In the presence of a magnetic field aligned in the $z$-direction with strength $B_{0}$, the nuclear spins of the chemical sample either align or anti-align with the direction of the field, leading to an energy difference between those aligned and anti-aligned given by $\Delta E=\gamma\hbar B_{0}$, where $\gamma$ is the gyromagnetic ratio of the 1H (or 13C) nuclei (values found at Ref. Stone, 2005). There are slightly more low energy, aligned spins than high energy, anti-aligned spins, giving the sample a bulk magnetization (vector) $M$ along the $z$-direction of the external field $B_{0}$. In FT-NMR, the sample is subjected to a radiofrequency (RF) pulse that perturbs the external magnetic field and creates a small, oscillating magnetic field in a direction perpendicular to $B_{0}$. The energy of this RF pulse is tuned to the field-dependent frequency, $f_{0}$, of the atomic nucleus one is targeting (e.g., 1H, 13C), via $f_{0}=\gamma B_{0}/2\pi$.Akitt (1983); Günther (1973) This perturbation changes the orientation of the nuclear spins in the sample, tipping the magnetization vector M away from the z-axis. After turning off the RF pulse, the nuclear spins relax back to their equilibrium configuration under the constant field, $B_{0}$. As these spins relax, the bulk magnetization $M$ also relaxes back to its equilibrium orientation along the $z$-axis, and the changing $M$ induces a voltage in the NMR detector coil. We call this signal the free induction decay (FID). The FT of the FID results in the familiar frequency domain spectrum that students learn in organic chemistry.
@@
The spectrum we obtain in 1H-NMR experiments provides insight into the molecular structure of a sample. After applying the FT to a temporal signal, a typical spectrum shows intensity as a function of frequency $\nu$ in units of Hertz. However, the convention is to convert this frequency in Hertz to a chemical shift ($\delta$) with units of parts-per-million (ppm) which reports on the change in frequency relative to the reference frequency of tetramethyl silane (TMS), $\nu_{\rm TMS}$, and the operational frequency of the spectrometer, $\nu_{\rm spec}$.Akitt (1983) Mathematically, $\delta=\tfrac{\nu-\nu_{\rm TMS}}{\nu_{\rm spec}}\times 10^{6}$. Importantly, this conversion yields a field-independent measurement and implies that the chemical shift does not change from instrument to instrument.
@@
Students are likely familiar with using features of a FT-NMR spectrum (e.g., chemical shifts, $J$-couplings, peak integrations), to characterize the structure of a molecule. It turns out that all of this same information is encoded into the FID, but in a more cryptic way (see Secs. II.1.1 and II.1.2). To see this, we recall that the FID signal encodes the precession frequency of each nucleus. Because this precession depends on the nuclei’s chemical and electrical environments, the detected signal is a sinusoid composed of frequencies unique to each nucleus’s environment. These signals combine to form the FID signal that the spectrometer measures at the detector. In addition, this FID signal decays to zero as the nuclear spins relax and approach equilibrium, with the relaxation being dictated by the chemical environment. The timescale of this relaxation is given by the spin-spin relaxation time ($T_{2}$), which helps determine the height and width of each peak in the spectrum.Akitt (1983) Because the FID signal is a combination of sinusoids with various frequencies and the signal decays to zero, it satisfies the requirements for the FT, as discussed in Sec. II.1.